OpenStudy (sgayton27):

Completely simplify: (cos θ − sin θ)^2

OpenStudy (campbell_st):

whats the expansion of the perfect square?

OpenStudy (campbell_st):

its just like expanding \[(a -b)^2=\]

OpenStudy (sgayton27):

Im sorry I was having trouble viewing your response but yes a^2 -2ab+b^2

OpenStudy (sgayton27):

I got cos^2x-2cosxsinx+sin^2x but it is not correct :/

OpenStudy (campbell_st):

you expansion is correct but what do you know about \[\cos^2(x) + \sin^2(x) \]

OpenStudy (campbell_st):

which are both terms in the expansion.

OpenStudy (sgayton27):

oh it equals 1!

OpenStudy (campbell_st):

so make the substitution and your answer should be correct